What is the next three terms of 3,5,9,15,23
Answer by greenestamps(13200)



This, like ANY OTHER similar question, can't be answered.

There is no way to know what the next term(s) are; ANY next three numbers will make a valid sequence.

You can look for an APPARENT pattern in the numbers to GUESS a POSSIBLE or even PROBABLE answer; but there is no way to know whether any answer you come up with is correct.

In this example, the difference between successive terms SUGGEST a POSSIBLE answer:

3 + 2 = 5
5 + 4 = 9
9 + 6 = 15
15 + 8 = 23

Since the differences are increasing by 2, the next three terms MIGHT BE

23 + 10 = 33
33 + 12 = 45
45 + 14 = 59

But, again.... That is only a GUESS.